# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 6. The Invariants of ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$
As an element of $S_{2}$, $u_{2}$ is a product of linear forms. Therefore $u_{2}$ divides $\mathcal{P}^{i}(u_{2})$ in $S_{2}$. Recall $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. For $0<i<q$, we have $\mathcal{P}^{i}(u_{2})\in R_{1}$. Since $u_{2}$ is linear in $\xi_{2}$, we have $\mathcal{P}^{i}(u_{2})=0$, proving (a). For $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$, we have $\mathcal{P}^{i}(u_{2})\in R_{2}$. Therefore $\mathcal{P}^{i}(u_{2})\in\ker(\Phi_{2,1})=u_{2}R_{2}$, proving (b). ∎
@@
(a) For $0<i<q$, $\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(d_{2,2})$ lie in $R_{2}$.
(b) For $0<i<q^{3}$, write $i=\ell q^{2}+r$ with $r<q^{2}$. Then $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{1,2})$ and $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{2,2})$ lie in $R_{3}$ with $\xi_{3}$-degree at most $\ell+1$.
@@
Suppose $0<i<q$. Using Lemma 6.5, $\mathcal{P}^{i}(u_{2})=0$. Therefore $\mathcal{P}^{i}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{i}(d_{2,2})$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$ and $i<q$, we have $\mathcal{P}^{i}(u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3})\in R_{2}$. Since $\mathcal{P}^{1}(\xi_{0}^{q}\xi_{3})=\xi_{0}^{q}\xi_{2}^{q}$ and $\mathcal{P}^{i}(\xi_{0}^{q}\xi_{3})=0$ for $1<i<q$, we see that $\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{2}$. Thus $u_{2}\mathcal{P}^{i}(d_{1,2})\in\ker(\Phi_{2,1})=u_{2}R_{2}$ and $\mathcal{P}^{i}(d_{1,2})\in R_{2}$. Similarly, since $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{2}$, proving $\mathcal{P}^{i}(d_{2,2})\in R_{2}$, completing the proof of (a).
The proof of (b) is by induction on $i$. Part (a) gives (b) for $i<q$. Suppose $i=\ell q^{2}+r$ with $r<q^{2}$ and $\ell<q$. Using Lemma 3.5, for $f\in R_{2}$ we have $\mathcal{P}^{i}(f)\in R_{3}$ with $\xi_{3}$-degree at most $\ell$. Since $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{3}$ with $\xi_{3}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. Similarly, using the observation $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$, we have $\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{3}$ with $\xi_{3}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. In the following, let $d$ denote either $d_{1,2}$ or $d_{2,2}$. Using the Cartan identity,
We have seen that $\mathcal{P}^{i}(u_{2}d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$. To prove (b), it is sufficient to show that for $j>0$, we have $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$. Write $i-j=\ell^{\prime}q^{2}+r^{\prime}$ with $r^{\prime}<q^{2}$ and $\ell^{\prime}\leq\ell$. By induction $u_{2}^{\ell^{\prime}+1}\mathcal{P}^{i}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell^{\prime}+1$.
Suppose $j<q^{2}$. Using Lemma 6.5, we have $\mathcal{P}^{j}(u_{2})/u_{2}\in R_{2}$. Therefore
$u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)=(\mathcal{P}^{i}(u_{2})/u_{2})u_{2}^{\ell+1}\mathcal{P}^{i-j}(d)\in R_{3}$
with $\xi_{3}$-degree at most $\ell+1$, as required.
For $q^{2}\leq j<q^{3}$, we have $\ell^{\prime}+1\leq\ell$ and $\mathcal{P}^{j}(u_{2})\in R_{3}$ with $\xi_{3}$-degree at most $1$. Therefore $u_{2}^{\ell}\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell$ and $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{3}$ with $\xi_{3}$-degree at most $\ell+1$, as required. ∎
@@
Suppose $i=(q-1)q^{n-2}+r$ with $r<q^{n-2}$ and $i\not\equiv_{(q)}0$. Then $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})$ has $\xi_{3}$-degree at most $q-1$
@@
From Lemma 6.6, we know that $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})\in R_{3}$ with with $\xi_{3}$-degree at most $q$. We will show that the $\xi_{3}$-degree at most $q-1$. Using the Cartan identity
Since $u_{2}^{q}d_{j,2}\in R_{3}$ with $\xi_{3}$-degree $1$, we see that $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})\in R_{3}$ with with $\xi_{3}$-degree at most $q$. However, by Lemma 6.4, the coefficient of $\xi_{3}$ in $u_{2}^{q}d_{j,2}$ is $u_{2}^{q-1}(u_{1}d_{j-1,1})^{q}$, which means that the coefficient of $\xi_{3}\xi_{2}^{q-1}$ in $u_{2}^{q}d_{j,2}$ is $(u_{1}^{q}d_{j-1,1})^{q}$. Therefore the coefficient of $\xi_{3}^{q}$ in $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})=0$ and the $\xi_{3}$-degree of $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is at most $q-1$. Since $i<q^{3}$ and $qk<i$, we have $k<q^{2}$. Thus $\mathcal{P}^{k}(u_{2})/u_{2}\in R_{2}$ and
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kp<(q-1)q^{2}$ and $u_{2}^{q}\mathcal{P}^{i-qk}(d_{j,2})$ has $\xi_{3}$-degree at most $q-1$. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$. ∎
This completes the proof of part (a) of Theorem 3.11 for $m=2$. Part (b) of Theorem 4.6 for $m=2$ follows from Lemmas 6.4 and 6.1. Part (c) of Theorem 4.6 is just Theorem 3.11.
The following gives part (d) of Theorem 4.6 for $m=2$.
@@
In $S_{2}^{{{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}}$, $\xi_{3}=(\xi_{2}+c_{2,2})d_{1,2}-\xi_{1}d_{2,2}-c_{3,2}$ with $c_{3,2}\in R_{2}$ and
$-c_{3,2}\equiv_{\langle\xi_{0}^{q+2}\rangle}\xi_{0}^{2}\xi_{1}^{q-2}\xi_{2}^{q-1}+\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}.$

### v3

@@ 6. The Invariants of $\text{O}^{+}_{{4}}(\mathbb{F}_{q})$}}
Suppose $0<i<q$. Using Lemma 6.4, $\mathcal{P}^{i}(u_{2})=0$. Therefore $\mathcal{P}^{i}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{i}(d_{2,2})$. Since $u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4}\in R_{3}$ and $i<q$, we have $\mathcal{P}^{i}(u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4})\in R_{3}$. Since $\mathcal{P}^{1}(\xi_{1}^{q}\xi_{4})=\xi_{1}^{q}\xi_{3}^{q}$ and $\mathcal{P}^{i}(\xi_{1}^{q}\xi_{4})=0$ for $1<i<q$, we see that $u_{2}\mathcal{P}^{i}(d_{1,2})=\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{3}$. It then follows from Lemma 6.3 that $\mathcal{P}^{i}(d_{1,2})\in R_{3}$. Similarly, since $u_{2}d_{2,2}-\xi_{2}^{q}\xi_{4}\in R_{3}$, we have $u_{2}\mathcal{P}^{i}(d_{2,2})=\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{3}$, proving $\mathcal{P}^{i}(d_{2,2})\in R_{3}$, completing the proof of (a).
The proof of (b) is by induction on $i$. Part (a) gives (b) for $i<q$. Suppose $i=\ell q^{2}+r$ with $r<q^{2}$ and $\ell<q$. Using Lemma 3.6, for $f\in R_{3}$ we have $\mathcal{P}^{i}(f)\in R_{4}$ with $\xi_{4}$-degree at most $\ell$. Using Lemma 6.6(a), $u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4}\in R_{3}$. Therefore $\mathcal{P}^{i}(u_{2}d_{1,2})\in R_{4}$ with $\xi_{4}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. Similarly, using Lemma 6.6(b), $u_{2}d_{2,2}-\xi_{2}^{q}\xi_{4}\in R_{3}$. Thus $\mathcal{P}^{i}(u_{2}d_{2,2})\in R_{4}$ with $\xi_{4}$-degree at most $1$ when $\ell=0$ and at most $\ell$ otherwise. In the following, let $d$ denote either $d_{1,2}$ or $d_{2,2}$. Using the Cartan identity,
We have seen that $\mathcal{P}^{i}(u_{2}d)\in R_{4}$ with $\xi_{4}$-degree at most $\ell+1$. To prove (b), it is sufficient to show that for $j>0$, we have $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{4}$ with $\xi_{4}$-degree at most $\ell+1$. Write $i-j=\ell^{\prime}q^{2}+r^{\prime}$ with $r^{\prime}<q^{2}$ and $\ell^{\prime}\leq\ell$. By induction $u_{2}^{\ell^{\prime}+1}\mathcal{P}^{i}(d)\in R_{4}$ with $\xi_{4}$-degree at most $\ell^{\prime}+1$.
Suppose $j<q^{2}$. Using Lemma 6.4, we have $\mathcal{P}^{j}(u_{2})/u_{2}\in R_{3}$. Therefore
$u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)=(\mathcal{P}^{j}(u_{2})/u_{2})u_{2}^{\ell+1}\mathcal{P}^{i-j}(d)\in R_{4}$
with $\xi_{4}$-degree at most $\ell+1$, as required.
For $q^{2}\leq j<q^{3}$, we have $\ell^{\prime}+1\leq\ell$ and $\mathcal{P}^{j}(u_{2})\in R_{4}$ with $\xi_{4}$-degree at most $1$. Therefore $u_{2}^{\ell}\mathcal{P}^{i-j}(d)\in R_{4}$ with $\xi_{4}$-degree at most $\ell$ and $u_{2}^{\ell}\mathcal{P}^{j}(u_{2})\mathcal{P}^{i-j}(d)\in R_{4}$ with $\xi_{4}$-degree at most $\ell+1$, as required. ∎
@@
Suppose $i=(q-1)q^{2}+r$ with $r<q^{2}$ and $i\not\equiv_{(q)}0$. Then $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})$ has $\xi_{4}$-degree at most q1q-1
@@
From Lemma 6.7, we know that $u_{2}^{q}\mathcal{P}^{i}(d_{j,2})\in R_{4}$ with $\xi_{4}$-degree at most $q$. We will show that the $\xi_{4}$-degree at most q1q-1. Using the Cartan identity
Since $u_{2}^{q}d_{j,2}\in R_{4}$ with $\xi_{4}$-degree $1$, we see that $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})\in R_{4}$ with $\xi_{4}$-degree at most $q$. However, by Lemma 6.6, the coefficient of $\xi_{4}$ in $u_{2}^{q}d_{j,2}$ is $u_{2}^{q-1}(u_{1}d_{j-1,1})^{q}$, which means that the coefficient of $\xi_{4}\xi_{3}^{q-1}$ in $u_{2}^{q}d_{j,2}$ is $(u_{1}^{q}d_{j-1,1})^{q}$. Therefore the coefficient of $\xi_{4}^{q}$ in $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})$. Since $r\not\equiv_{(q)}0$, $\mathcal{P}^{r}((u_{1}^{q}d_{j-1,1})^{q})=0$ and the $\xi_{4}$-degree of $\mathcal{P}^{i}(u_{2}^{q}d_{j,2})$ is at most q1q-1. Since $i<q^{3}$ and $qk<i$, we have $k<q^{2}$. Thus $\mathcal{P}^{k}(u_{2})/u_{2}\in R_{3}$ and
Write $r=r_{0}+qr^{\prime}$ with $0<r_{0}<q$. If $r^{\prime}=0$, then $i-kq<(q-1)q^{2}$ and $u_{2}^{q}\mathcal{P}^{i-qk}(d_{j,2})$ has $\xi_{4}$-degree at most q1q-1. If $r^{\prime}>0$, then the result follows by induction on $r^{\prime}$. ∎
@@
In $S_{2}^{{{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}}$, $\xi_{4}=(\xi_{3}+c_{3,2})d_{1,2}-\xi_{2}d_{2,2}-c_{4,2}$ with $c_{4,2}\in R_{3}$ and
$c_{4,2}\equiv_{\langle\xi_{1}^{q+2}\rangle}\xi_{1}^{2}\xi_{2}^{q-2}\xi_{3}^{q-1}+\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3}.$
@@
Using Lemma 6.1, $u_{2}=\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1}$ with $c_{3,2}\in R_{2}$. Since $u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4}$ and $u_{2}d_{2,2}-\xi_{2}^{q}\xi_{4}$ are elements of $R_{3}$ (see Lemma 6.6),
$u_{2}\xi_{4}-(\xi_{3}+c_{3,2})u_{2}d_{1,2}+\xi_{2}u_{2}d_{2,2}\in R_{3}.$
Since this expression is divisible by $u_{2}$ in $S_{2}$ and lies in $R_{3}$, it is divisible by $u_{2}$ in $R_{3}$ (see Lemma 6.3). Defining $c_{4,2}:=(u_{2}\xi_{4}-(\xi_{3}+c_{3,2})u_{2}d_{1,2}+\xi_{2}u_{2}d_{2,2})/u_{2}$ gives
$\xi_{4}=(\xi_{3}+c_{3,2})d_{1,2}-\xi_{2}d_{2,2}-c_{4,2}\ .$
with $c_{4,2}\in R_{3}$. Using Lemma 6.6
$(\xi_{3}+c_{3,2})u_{2}d_{1,2}-\xi_{2}u_{2}d_{2,2}$ $\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $\xi_{4}((\xi_{3}+c_{3,2})\xi_{1}^{q}-\xi_{2}^{q+1})-\xi_{3}(\xi_{2}\xi_{3}^{q}-\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1})$
$+\xi_{2}(\xi_{3}^{q+1}+\xi_{1}^{q}\xi_{2}^{q^{2}-q}\xi_{3})-c_{3,2}(\xi_{2}\xi_{3}^{q}-\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1})$
$=$ $u_{2}\xi_{4}+\xi_{1}\xi_{2}^{q}\xi_{3}^{q}+\xi_{1}^{q}\xi_{2}^{q^{2}-q+1}\xi_{3}$
$-c_{3,2}(\xi_{2}\xi_{3}^{q}-\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1})$
Therefore, using Lemma 6.1 to substitute for $c_{3,2}$,
$-u_{2}c_{4,2}$ $\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $-\xi_{1}\xi_{2}^{q}\xi_{3}^{q}-\xi_{1}^{q}\xi_{2}^{q^{2}-q+1}\xi_{3}+c_{3,2}(\xi_{2}\xi_{3}^{q}-\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1})$
$\equiv_{\langle\xi_{1}^{q+2}\rangle}$ $-\xi_{1}\xi_{2}^{q}\xi_{3}^{q}-\xi_{1}^{q}\xi_{2}^{q^{2}-q+1}\xi_{3}+\xi_{1}\xi_{2}^{q-1}(\xi_{2}\xi_{3}^{q}-\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1})$
$=$ $-\xi_{1}^{2}\xi_{2}^{2q-1}\xi_{3}^{q-1}-\xi_{1}^{q}\xi_{2}^{q^{2}-q+1}\xi_{3}$
$=$ $-\xi_{2}^{q+1}(\xi_{1}^{2}\xi_{2}^{q-2}\xi_{3}^{q-1}+\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3}).$
Thus $c_{4,2}\equiv_{\langle\xi_{1}^{q+2}\rangle}\xi_{1}^{2}\xi_{2}^{q-2}\xi_{3}^{q-1}+\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3}.$
Applying $\mathcal{P}^{1}$ to $\xi_{4}=(\xi_{3}+c_{3,2})d_{1,2}-\xi_{2}d_{2,2}-c_{4,2}$ and using $\mathcal{P}^{1}(c_{3,2})=\xi_{2}^{q}$ from Lemma 6.5 gives
$\xi_{3}^{q}=2\xi_{2}^{q}d_{1,2}-2\xi_{1}^{q}d_{2,2}-\mathcal{P}^{1}(c_{4,2})+(\xi_{3}+c_{3,2})\mathcal{P}^{1}(d_{1,2})-\xi_{2}\mathcal{P}^{1}(d_{2,2}).$
Since $c_{4,2}\in R_{3}$, $\mathcal{P}^{1}(c_{4,2})\in R_{3}$. Furthermore, using Lemma 6.9,
$-\mathcal{P}^{1}(c_{4,2})$ $\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $\mathcal{P}^{1}(\xi_{1}^{2}\xi_{2}^{q-2}\xi_{3}^{q-1}+\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3})$
$\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $2\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}-\xi_{1}^{2}\xi_{2}^{2q-2}\xi_{3}^{q-2}+\xi_{1}^{q}\xi_{2}^{q^{2}-q}\ .$
Using Lemmas 6.3, 6.4, 6.6 and 6.7, $\mathcal{P}^{1}(u_{2}d_{1,2})=u_{2}\mathcal{P}^{1}(d_{1,2})$, $\mathcal{P}^{1}(d_{1,2})\in R_{3}$ and
$u_{2}\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}^{q}-2\xi_{1}^{q}\xi_{3}^{q}+\xi_{2}^{q+1}\xi_{3}^{q-1}-\xi_{1}\xi_{2}^{2q}\xi_{3}^{q-2}=\xi_{2}^{q+1}\xi_{3}^{q-1}-\xi_{1}\xi_{2}^{2q}\xi_{3}^{q-2}-\xi_{1}^{q}\xi_{3}^{q}.$
Again using Lemmas 6.3, 6.4, 6.6 and 6.7, $\mathcal{P}^{1}(u_{2}d_{2,2})=u_{2}\mathcal{P}^{1}(d_{2,2})$, $\mathcal{P}^{1}(d_{2,2})\in R_{3}$ and $u_{2}\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\xi_{3}^{q}-\xi_{2}^{q}\xi_{3}^{q}-\xi_{1}^{q}\xi_{2}^{q^{2}}=-\xi_{1}^{q}\xi_{2}^{q^{2}}$. Since $u_{2}\equiv_{\langle\xi_{1}^{q+1}\rangle}-\xi_{2}^{q+1}+\xi_{1}^{q}\xi_{3}$, we have
$\mathcal{P}^{1}(d_{1,2})\equiv_{\langle\xi_{1}^{q+1}\rangle}-\xi_{3}^{q-1}+\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-2}$
and $\mathcal{P}^{1}(d_{2,2})\equiv_{\langle\xi_{1}^{q+1}\rangle}\xi_{1}^{q}\xi_{2}^{q^{2}-q-1}$. Since $\mathcal{P}^{1}(d_{1,2})$ and $\mathcal{P}^{1}(d_{2,2})$ lie in $R_{3}$, we have $\xi_{2}^{q}d_{1,2}-\xi_{1}^{q}d_{2,2}\in R_{3}$. Using the congruences we have
$\xi_{3}^{q}$ $\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $2\xi_{2}^{q}d_{1,2}-2\xi_{1}^{q}d_{2,2}+(2\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}-\xi_{1}^{2}\xi_{2}^{2q-2}\xi_{3}^{q-2}+\xi_{1}^{q}\xi_{2}^{q^{2}-q})$
$-(\xi_{3}+c_{3,2})(\xi_{3}^{q-1}-\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-2})-\xi_{1}^{q}\xi_{2}^{q^{2}-q}$
$\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $2\xi_{2}^{q}d_{1,2}-2\xi_{1}^{q}d_{2,2}+3\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}-\xi_{3}^{q}-\xi_{1}^{2}\xi_{2}^{2q-2}\xi_{3}^{q-2}$
$-c_{3,2}(\xi_{3}^{q-1}-\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-2})$
$\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $2\xi_{2}^{q}d_{1,2}-2\xi_{1}^{q}d_{2,2}-\xi_{3}^{q}+3\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}-\xi_{1}^{2}\xi_{2}^{2q-2}\xi_{3}^{q-2}$
$-\xi_{1}\xi_{2}^{q-1}(\xi_{3}^{q-1}-\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-2})$
$\equiv_{\langle\xi_{1}^{q+1}\rangle}$ $2\xi_{2}^{q}d_{1,2}-2\xi_{1}^{q}d_{2,2}-\xi_{3}^{q}+2\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}.$
Therefore
$\xi_{3}^{q}\equiv_{\langle\xi_{1}^{q+1}\rangle}\xi_{2}^{q}d_{1,2}-\xi_{1}^{q}d_{2,2}+\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}.$
Recall that $\xi_{2}^{q}d_{1,2}+\xi_{1}^{q}d_{2,2}\in R_{3}$ and define $r_{2,1}\in R_{3}$ by
(5) $\xi_{1}^{q+1}r_{2,1}=\xi_{3}^{q}-\xi_{2}^{q}d_{1,2}+\xi_{1}^{q}d_{2,2}-\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}.$
Then
(6) $\xi_{3}^{q}=\xi_{2}^{q}d_{1,2}-\xi_{1}^{q}d_{2,2}-\gamma_{3,2}^{(1)}$
with $\gamma_{3,2}^{(1)}=-\xi_{1}^{q+1}r_{2,1}-\xi_{1}\xi_{2}^{q-1}\xi_{3}^{q-1}\in R_{3}$ and $\nu(\gamma_{3,2}^{(1)})>q+1$.
We have completed the proof of Lemma 4.6 for the base case $m=2$. Parts (a), (c) and (d) follow from Lemmas 6.1 and 6.6. Part (b) follows from Lemmas 6.4, 6.7 and 6.8. Part (e) follows from Lemma 6.9 and part (f) is Equation 6. Theorem 4.8 for $m=2$, gives the alternative computation for the invariants of $\text{O}^{+}_{{4}}(\mathbb{F}_{q})$}}.
Remark 6.10.
Notice that in the process of deriving Equation (6), the expression $\xi_{3}^{q}$ arises on the right hand side. If the coefficient had been +1 rather than -1 then this process would not have provided a formula for $\xi_{3}^{q}$. It is for this reason that naively applying a Steenrod operation to the formula for $\xi_{n-i}^{q^{i}}$ is not guaranteed to give a formula for $\xi_{n-1-i}^{q^{i+1}}$.
7. Proof of Lemma 4.6
The proof is by induction on $m$. The base case, $m=2$, is given in Section 6. For the induction step, we will assume the lemma is true for $S_{\ell}^{G_{\ell}}$ for $2\leq\ell<m$ and, in the proof of a given part, we will assume the earlier parts are true for $S_{m}^{G_{m}}$.
@@
Part (e) of Lemma 4.6 follows from parts (a) and (c).